# Open-Vocabulary SAM3D: Understand Any 3D Scene

### v2

@@ 4 Experiments
To quantitatively evaluate our method, we compare OV-SAM3D with some leading open-vocabulary 3D scene understanding methods and 3D instance segmentation methods suitable for 3D scene understanding tasks on the well-studied dataset ScanNet200 [31] and the outdoor dataset nuScenes [5]. Furthermore, we provide ablation studies for OV-SAM3D, explaining the improvement of each module in our framework and the impact of different hyper-parameters. In addition, we display some qualitative results of our method in any 3D scene, more qualitative results can be seen in Appendix. B.
@@ 4.1 Experimental Setup
ScanNet200 [31] is a RGB-D dataset frequently utilized for evaluating open-vocabulary 3D scene understanding tasks, which expands a richer 200 instance categories based on the basic ScanNet dataset [9]. ScanNet200 is composed of 1513 indoor scenes divided into 1201 training scenes and 312 validation scenes, which contains surface mesh point cloud, instance-level semantic annotation of scenes, and semantic labels of each point. And nuScenes [5] is an outdoor dataset including 850 training scenes, 150 validation scenes, and 150 testing scenes, which contains various sensor data to achieve multiple 3D perception tasks such as 3D object detection, instance segmentation, panoramic segmentation, object tracking. Considering that existing open-vocabulary 3D scene understanding algorithms have hardly known the outdoor dataset, we conduct some supplementary experiments on nuScenes to verify the effectiveness of OV-SAM3D for any scene.
@@
Following the setting of the baseline [36], we choose the average precision score to evaluate the quantitative results. We respectively calculate the score at the average mask overlap thresholds between 50% and 95% along 5% steps as AP and at the mask overlap threshold of 25%, 50% as $\mathrm{AP_{25}}$, $\mathrm{AP_{50}}$. For ScanNet200 dataset, we separately calculate AP under different category groups [31] as $\mathrm{AP_{head}}$, $\mathrm{AP_{com}}$ and $\mathrm{AP_{tail}}$. For nuScenes dataset, we separately calculate the average precision for the three main categories of car, truck, and bus as $\mathrm{AP_{car}}$, $\mathrm{AP_{truck}}$ and $\mathrm{AP_{bus}}$.
@@
As an open-vocabulary 3D scene understanding framework that does not require any prior knowledge of datasets, OV-SAM3D sets some interactive hyper-parameters for arbitrary 3D scene. In order to compare with other existing algorithms on the close-dataset, we can only choose to fix these hyper-parameters, which strongly affect the performance of our algorithm. For ScanNet200, we select 200 high-quality superpoints as initial prompts, set the threshold $\tau$ in Eq. 1 to 0.45 and sample RGB-D frames at a frequency of 10. For nuScenes, we select 100 high-quality superpoints and keep the threshold $\tau$. In addition, more details about converting nuScenes data formats shown in Appendix. A. We complete our experiment on a single GPU.
@@ 4.2 Quantitative 3D scene understanding evaluation
For the purpose of maintaining the same experimental conditions as other existing open-vocabulary 3D scene understanding methods, we conduct a quantitative experiment on the well-studied dataset ScanNet200. However, some methods actually pre-learned the relevant knowledge about ScanNet200, which is not in line with our original intention of understanding any 3D scene in unknown environments. Subsequently, we conduct further experiments on the nuScenes dataset to demonstrate the advantages of our method in understanding any 3D scene.
@@
The numerical results on ScanNet200 are presented in Tab. 1, where we compare OV-SAM3D with two types of algorithms: one that requires pre-training on ScanNet200, and the other that can be adapted to open-vocabulary 3D scene understanding tasks without training. In the environment that allows for learning in advance, our advantage of understanding any open-world 3D scene seems difficult to demonstrate, but even so, we reflect a certain level of competitiveness especially in $\mathrm{AP_{25}}$ and $\mathrm{AP_{tail}}$. In order to further highlight our research objectives, we compare with the state-of-the-art method OpenMask3D on the outdoor nuScenes dataset that are blank for all algorithms.
Table 1: 3D scene understanding results on the ScanNet200 validation set. In the listed heading methods, the best results are in bold while the second best results are underscored. We compare against both methods with training and methods without training, and report different evaluation indicators, in which $\mathrm{AP_{tail}}$ can evaluate the performance of models on the long-tail distribution.
@@
As a high-quality dataset for autonomous driving, nuScenes is rarely used in the task of open-vocabulary 3D scene understanding. Therefore, we supplement some experiment on such an unseen dataset. As shown in Tab. 2, the state-of-the-art method OpenMask3D with published weights claiming applicable to arbitrary scene is clearly not competent on the outdoor nuScenes dataset, while our OV-SAM3D still maintains impressive understanding performance.
Table 2: 3D scene understanding results on the nuScenes validation set.
Method Setting 3D Proposal Semantic $\mathrm{AP}$ $\mathrm{AP_{50}}$ $\mathrm{AP_{25}}$ $\mathrm{AP_{car}}$ $\mathrm{AP_{truck}}$ $\mathrm{AP_{bus}}$
OpenMask3D [36] Specific Training Mask3D [35] CLIP [27] 0.5 1.4 5.2 0.3 0.2 0.6
Ours Without Training None CLIP [27] 8.9 16.0 29.1 13.6 16.2 39.8
@@ 4.3 Qualitative results
Here we share some qualitative results from our OV-SAM3D for understanding any 3D scene. In Fig. 6, our approach achieves the 3D scene understanding in an open-world environment, automatically recognizing various instances in the scene without the need for text query with its embedded open-tags. In Fig. 7, we present the text-driven results of OV-SAM3D on the ScanNet200 dataset and nuScenes dataset. Not limited to simple category text queries, OV-SAM3D demonstrates the ability to further understand the entire 3D scene and detect objects under the guidance of understanding complex languages.
@@ 4.4 Ablation study
In this section, we discuss the influence of different components of OV-SAM3D on scene understanding, and the effectiveness of each module design can be proved from Tab. 3. As mentioned in the Sec. 4.1, OV-SAM3D retains some interactive hyper-parameters, and in Tab.4, we present the results of different hyper-parameters.
@@
The interactive hyper-parameters in OV-SAM3D are designed for users to understand any 3D scene. On large-scale datasets, such fixed hype-rparameters are difficult to accommodate all scenes. In Tab. 4, we demonstrate the potential impact of different hyper-parameter selections.
Table 4: Ablation studies on hyper-parameters. The experiments are conducted on ScanNet200 validation set and we keep selecting tags with the highest feature similarity as instance labels.
@@
Figure 7: Qualitative results of our methods on the ScanNet200 and nuScenes datasets with text queries.
@@ 5 Conclusion and Future Works
We propose OV-SAM3D, a pioneering framework that can achieve open-vocabulary 3D scene understanding tasks in an open-world environment. Compared to existing open-vocabulary 3D scene understanding algorithms can achieve the understanding tasks on arbitrary scene without the need to learn any prior knowledge about the given scene. OV-SAM3D sufficiently exploits the generalization ability of the foundation models, among which the designed overlapping score can effectively utilize the powerful zero-shot segmentation ability of SAM, and the critical open image tags come from the RAM. Based on the cooperation of foundation models, we expand OV-SAM3D to a more challenging and meaningful open-world environment, and this attempt in the open-world has also been advocated in the related research of foundation models. Comparative experiments on the ScanNet200 and nuScenes datasets indicate that OV-SAM3D not only achieves competitive results in the setting of well-studied ScanNet200 dataset, but also achieves a significant lead in open nuScenes dataset, which demonstrates that OV-SAM3D is capable of understanding any 3D scene in an unknown open-world.
Limitations.
As a universal open-vocabulary 3D scene understanding model, the understanding ability of OV-SAM3D comes entirely from the foundation model. Therefore, OV-SAM3D cannot understand some complex 3D scenes well derived from this fact the vision foundation models cannot handle overly complex tasks with zero-shot. However, while foundational models limit our current capabilities, they also indicate the direction for our future work. We believe that with the growth of foundation model capabilities, our OV-SAM3D will gradually make significant progress.
@@
  • Sec. A shows more experimental processing details for nuScenes dataset.
  • Sec. B shows more visualization results of Open Vocabulary SAM3D.
@@ Appendix A Implementation Details on nuScenes
The nuScenes dataset provides comprehensive autonomous driving data collected in various outdoor environments. In contrast, the ScanNet200 dataset focuses on indoor scenes, resulting in different data representations and types. To achieve a unified evaluation framework, it is crucial to preprocess the nuScenes data, particularly its point clouds and instance annotations, to conform to the format and structure of ScanNet200. This process involves two main tasks: point cloud processing and instance assignment.
@@ A.1 Point Cloud Processing
In ScanNet200, point cloud data is typically stored in a mesh format, directly providing rich surface information. However, nuScenes stores point clouds as raw coordinate data. To ensure the generation of superpoints, we convert the nuScenes point clouds into mesh representations.
@@
The conversion from raw point clouds to meshes involves surface reconstruction techniques. To achieve this, we employ an Alpha-Shape-based surface reconstruction method. This approach involves point cloud preprocessing, Alpha shape computation, mesh generation, and optimization to construct a polygon mesh from the point cloud data and restore the surface structure. To balance noise handling and detail preservation, we define the parameter $\mathrm{\alpha=0.5}$, ultimately converting the nuScenes point cloud data into mesh representations compatible with the ScanNet200 data structure.
@@
Superpoints are generated based on the mesh, and we ensure that the parameters matched those used in ScanNet200. Specifically, we select the top 100 superpoints based on the number of points in the cloud to maintain high relevance and detail.
@@ A.2 Instance Assignment
nuScenes lacks explicit point cloud instance annotations, which are crucial for various qualitative evaluations. To address this issue, we utilize the bounding box annotations provided by nuScenes to infer instance-level segmentation.
@@
The instance assignment process involves segmenting the point cloud based on bounding box annotations. We extract the point cloud fragments that fall within each annotated box, effectively creating instance-level segments.
@@
Isolate points within each annotated bounding box to form preliminary instance fragments. To ensure that each instance fragment has sufficient point cloud density, we set the point cloud for each instance to $\mathrm{P_{points\_per\_sample}>100}$.
@@
nuScenes provides a rich set of annotated box categories. To balance the number of instances and point cloud density, we filtered and merged the annotation box categories, ultimately obtaining ten common category labels.

### v3

@@ 4 Experiments
To quantitatively evaluate our method, we compare OV-SAM3D with some leading open-vocabulary 3D scene understanding methods and 3D instance segmentation methods suitable for 3D scene understanding tasks on the well-studied dataset ScanNet200 [33] and the outdoor dataset nuScenes [4]. Furthermore, we provide ablation studies for OV-SAM3D, explaining the improvement of each module in our framework and the impact of different hyper-parameters. In addition, we display some qualitative results of our method in any 3D scene, more qualitative results can be seen in the Appendix.
@@ 4.1 Experimental Setup
ScanNet200 [33] is an RGB-D dataset commonly utilized for evaluating open-vocabulary 3D scene understanding tasks, which expands the original ScanNet dataset [8] by labeling a more extensive set of 200 instance categories. ScanNet200 consists of 1,513 indoor scenes divided into 1,201 training scenes and 312 validation scenes, which contains surface mesh point clouds, instance-level semantic annotations, and semantic labels for each point. In contrast, nuScenes [4] is an outdoor dataset comprising 850 training scenes, 150 validation scenes, and 150 testing scenes. It provides various sensor data for multiple 3D perception tasks, such as 3D object detection, instance segmentation, panoramic segmentation, and object tracking. Given that existing open-vocabulary 3D scene understanding algorithms are rarely applied to outdoor datasets, we conduct supplementary experiments on nuScenes to evaluate the effectiveness of OV-SAM3D across diverse scenes.
@@
Following the setting of the baseline [37], we choose the average precision score to evaluate the quantitative results. We respectively calculate the score at the average mask overlap thresholds between 50% and 95% along 5% steps as AP and at the mask overlap threshold of 25%, 50% as $\mathrm{AP_{25}}$, $\mathrm{AP_{50}}$. For ScanNet200 dataset, we calculate AP under different category groups [33], namely $\mathrm{AP_{head}}$, $\mathrm{AP_{com}}$ and $\mathrm{AP_{tail}}$. For the nuScenes dataset, we calculate the average precision for the three main categories of car, truck, and bus as $\mathrm{AP_{car}}$, $\mathrm{AP_{truck}}$ and $\mathrm{AP_{bus}}$.
@@
As an open-vocabulary 3D scene understanding framework that does not require prior knowledge of datasets, OV-SAM3D sets some interactive hyper-parameters for arbitrary 3D scenes. To enable comparison with other existing algorithms on the close-dataset, we fix these hyper-parameters, which significantly affects the performance of our algorithm. For ScanNet200, we select 200 high-quality superpoints as initial prompts, set the threshold $\tau$ in Eq. 1 to 0.45 and sample RGB-D frames at a frequency of 10. For nuScenes, we select 100 high-quality superpoints and keep the threshold $\tau$. In addition, more details about converting nuScenes data formats are provided in the Appendix. Following these settings, the task of performing a comprehensive scene understanding, which includes both segmentation and recognition, requires an average of 4-6 minutes on a ScanNet scene and 30-60 seconds on a nuScenes scene when executed on a single GPU.
@@ 4.2 Quantitative 3D scene understanding evaluation
To ensure consistency with existing open-vocabulary 3D scene understanding methods, we conducted quantitative experiments on the well-studied ScanNet200 dataset. However, it is important to note that some methods had prior knowledge of ScanNet200, which diverges from our original goal of understanding any 3D scene in unknown environments. Therefore, we extended our experiments to the nuScenes dataset to further demonstrate the superiority of our method in understanding any 3D scene.
@@
The numerical results on ScanNet200 are presented in Tab. 1, where we compare OV-SAM3D with two types of algorithms: those requiring pre-training on ScanNet200 and those adaptable to open-vocabulary 3D scene understanding tasks without such training. In scenarios where pre-training is permitted, our advantage of understanding any open-world 3D scene seems difficult to demonstrate, but even so, we reflect a certain level of competitiveness especially in $\mathrm{AP_{25}}$ and $\mathrm{AP_{tail}}$. In order to further highlight our research objectives, we compare with the state-of-the-art method OpenMask3D on the outdoor nuScenes dataset that is blank for all algorithms.
@@
Table 1: 3D scene understanding results on the ScanNet200 validation set. In the table, the best results are in bold while the second best results are underscored. We compare against both methods with training and methods without training, and report different evaluation indicators, in which $\mathrm{AP_{tail}}$ can evaluate the performance of models on the long-tail distribution.
@@
As a high-quality dataset for autonomous driving, the nuScenes dataset is seldom utilized in the task of open-vocabulary 3D scene understanding. Therefore, we supplement some experiment on this less-explored dataset to compare the capabilities of models in the open world. Most advanced methods that require training [27, 24] have been trained exclusively on indoor datasets, and directly applying or finetuning them on the nuScenes dataset would undermine the validity of the experiment. Therefore, we select OpenMask3D [37], which provides a publicly available version capable of running on arbitrary scenes. For methods that do not require training, the typical approach involves a two-stage process: first, generating class-agnostic 3D masks, followed by integration with existing 3D scene understanding frameworks. In our experiments, we consistently utilized OpenMask3D for this phase. As shown in Tab. 2, these state-of-the-art methods are clearly not competent on the outdoor nuScenes dataset, while our OV-SAM3D still maintains impressive understanding performance.
Method Setting $\mathrm{AP}$ $\mathrm{AP_{50}}$ $\mathrm{AP_{25}}$ $\mathrm{AP_{car}}$ $\mathrm{AP_{truck}}$ $\mathrm{AP_{bus}}$
OpenMask3D [37] Training for arbitrary scenes 0.5 1.4 5.2 0.3 0.2 0.6
SAMPro3D [41] + OpenMask3D Training-free 0.8 1.1 5.2 0.6 0.0 0.2
SAI3D [43] + OpenMask3D Training-free 2.8 5.4 7.2 1.9 2.6 22.4
Ours Training-free 8.9 16.0 29.1 13.6 16.2 39.8
Table 2: 3D scene understanding results on the nuScenes validation set. Following the setup of SAI3D, we uniformly use OpenMask3D as the model for matching class-agnostic masks with text features. For the complete OpenMask3D, we choose its publicly available model that adapts to any 3D scene.
@@ 4.3 Qualitative results
Here we show some qualitative results from our OV-SAM3D. In Fig. 6, our approach achieves the 3D scene understanding in an open-world environment, automatically recognizing various instances in the scene without requiring a text query, due to its embedded open tags. In Fig. 7, we present the text-driven results of OV-SAM3D on the ScanNet200 dataset and nuScenes dataset. Not limited to simple category text queries, OV-SAM3D demonstrates an enhanced ability to comprehend the entire 3D scene and detect objects, even with complex language cues.
@@
Figure 7: Qualitative results of our method on the ScanNet200 datasets and the nuScenes datasets with various text queries.
@@ 4.4 Ablation study
In this section, we discuss the influence of different components of OV-SAM3D on scene understanding, and the effectiveness of each module design can be proved from Tab. 3. As mentioned before, OV-SAM3D retains some interactive hyper-parameters, and in Tab.4, we present the results of different hyper-parameters.
@@
The interactive hyper-parameters in OV-SAM3D are designed for users to understand any 3D scene. On large-scale datasets, such fixed hyperparameters are difficult to accommodate all scenes. In Tab. 4, we demonstrate the potential impact of different hyper-parameter selections.
@@
Table 4: Ablation studies on hyperparameters. The experiments are conducted on ScanNet200 validation set and we keep selecting tags with the highest feature similarity as instance labels.
@@ 5 Conclusion and Future Works
We propose OV-SAM3D, a pioneering framework for open-vocabulary 3D scene understanding in open-world environments. Compared to existing algorithms, OV-SAM3D can understand arbitrary scenes without prior knowledge of the given environment. Through effectively exploring and leveraging the complementary strengths of foundational models, we have extended OV-SAM3D to a more challenging and meaningful open-world environment, a direction that has also been advocated in related foundational model research. Comparative experiments on the ScanNet200 and nuScenes datasets reveal that OV-SAM3D not only delivers competitive results on the well-studied ScanNet200 dataset but also significantly outperforms on the open nuScenes dataset, demonstrating its capability to understand any 3D scene in an unknown open-world. As a universal open-vocabulary 3D scene understanding model, OV-SAM3D inherits its comprehension capabilities from the foundational models, which is both its limitation and advantage. With the gradual development of foundation models, OV-SAM3D is expected to achieve further significant breakthroughs.
@@
  • Sec.A A shows more experimental processing details for nuScenes dataset.
  • Sec.B B shows more visualization results of Open Vocabulary SAM3D.
@@ Appendix A Implementation Details on nuScenes
As a commonly used dataset for 3D segmentation and understanding, the preprocessing and evaluation procedures of the ScanNet200 dataset are already quite mature, and we follow the same setup as previous work. Compared to the ScanNet200 dataset, which focuses on indoor scenes, the nuScenes dataset provides comprehensive autonomous driving data collected in various outdoor environments, resulting in different data representations and types. To achieve a unified evaluation framework, it is crucial to preprocess the nuScenes data, particularly its point clouds and instance annotations, to conform to the format and structure of ScanNet200. This process involves two main tasks: point cloud processing and instance assignment.
@@ A.1 Point Cloud Processing
In the ScanNet200 dataset, point cloud data is typically stored in a mesh format, directly providing detailed surface information. However, the nuScenes dataset stores point clouds as raw coordinate data. To facilitate the generation of superpoints, we should convert the nuScenes point clouds into mesh representations first.
@@
The conversion from raw point clouds to meshes involves employing surface reconstruction techniques. Specifically, we use an Alpha-Shape-based surface reconstruction method, which consists of several steps: point cloud preprocessing, Alpha shape computation, mesh generation, and optimization. These steps are used to create a polygon mesh from the point cloud data, effectively restoring the surface structure. To balance noise reduction and detail preservation, we set the parameter $\alpha$ to 0.5, ultimately converting the nuScenes point cloud data into mesh representations compatible with the ScanNet200 data structure.
@@
Superpoints are generated based on the mesh, ensuring that the parameters align with those used in the ScanNet200 dataset. Specifically, we select the top 100 superpoints as the initial 3D prompts based on the number of points in the cloud to maintain high relevance and detail.
@@ A.2 Instance Assignment
The nuScenes dataset lacks explicit point cloud instance annotations, which are critical for various qualitative evaluations. To overcome this limitation, we employ the bounding box annotations provided by nuScenes to infer instance-level segmentation.
@@
The instance assignment process involves segmenting the point cloud based on bounding box annotations. We extract the point cloud fragments within each annotated box, effectively creating instance-level segments.
@@
The instance assignment process entails segmenting the point cloud based on bounding box annotations. We isolate the point cloud fragments within each annotated bounding box, thereby creating preliminary instance fragments. To ensure each segment has sufficient point cloud density, we filter out annotation instances containing an insufficient number of points.
@@
The nuScenes dataset offers a comprehensive set of annotated box categories. To balance the number of instances and ensure sufficient point cloud density, we filtered and merged the annotation box categories, resulting in ten common category labels.